OpenStudy (anonymous):

solve the equation 2tanC-3=3 tan C-4 algebraically for all values of C in the interval 0

OpenStudy (mathmale):

Let me see whether I'm reading your problem statement correctly. Is this right?\[2*\tan(c-3) = 3*\tan (c-4)\]

OpenStudy (anonymous):

\[2 \tan C-3=3 \tan C-4\] thts exactly the way its written in my paper

OpenStudy (anonymous):

\[0 \le C < 360\]

OpenStudy (anonymous):

\[0 \le C < 360\]

OpenStudy (anonymous):

hello?

OpenStudy (mathmale):

Knowing this helps a LOT. Where you have \(2 \tan C-3=3 \tan C-4\) please substitute x for "tan C" because this simplifies the solution of the problem. Mind writing that out?

OpenStudy (anonymous):

I got 2tan c=3tan c-1

OpenStudy (mathmale):

I'd just made a suggestion that I'd thought might simplify the problem solution for you. If you take my suggestion, you'd get 2x=3x-1. Would you mind solving that for x?

OpenStudy (mathmale):

(Remember that this x represents tan C.)

OpenStudy (mathmale):

If 2x=3x-1, please add 1 to both sides and simplify the result.

OpenStudy (anonymous):

oh so I got -tan c=-1?

OpenStudy (anonymous):

or -1=x

OpenStudy (anonymous):

?

OpenStudy (mathmale):

My result is x=1. Would you please check your work? However, your tan C = 1 is correct. Are you able to visualize which angles from [0,360] have the tangent equal to 1?

OpenStudy (anonymous):

2tanC-3tanC= -tanC??

OpenStudy (anonymous):

nvm got it

OpenStudy (mathmale):

What you are doing is evaluating \(\tan ^{-1}1,\), or, in other woerds, you're finding the angle that has a tangent of 1.

OpenStudy (anonymous):

tan1=45?

OpenStudy (mathmale):

Which two angles in [0,360] have a tangent = to 1? Of course, Lyubas, you don't have to accept or use my suggestion of replacing tan C with x temporarily. The angle whose tangent is 1 is 45 degrees, yes, but there is another angle. Can you find it?

OpenStudy (mathmale):

Please type this: tan1=45? as either arctan 1 = 45 degrees, or \[\tan ^{1-}1=45 \deg\]

OpenStudy (anonymous):

I got tht so how do I find the 2nd angle?

OpenStudy (anonymous):

or is this it?

OpenStudy (mathmale):

Here's what to look for: the tangent function is defined as \[\tan \theta=\frac{ opp }{ adj }\] where opp=opposite side and adj=adjacent side. OK with that?

OpenStudy (anonymous):

yes

OpenStudy (mathmale):

note that tan 45 deg = 1/1 (both the opp side and the adj side are 1). Note that tan 225 deg = -1/(-1) (both the opp side and the adj side are equal to -1 in the third quadrant). So the 2nd solution is 225 deg. You might want to draw this situation; doing so may make the situation clearer for you.

OpenStudy (anonymous):

got it thnks answer : 45,225 ?

OpenStudy (mathmale):

Yes! You could improve this answer by writing it as theta: {45, 225}.

OpenStudy (mathmale):

If you're completely satisfied with this solution and the methods we used, move on to another problem. I'm sorry some of my responses have been slow; I have several others also waiting for help.

OpenStudy (anonymous):

alrighty thx

OpenStudy (mathmale):

Great working with you!

OpenStudy (anonymous):

same here:)